Circular No. 5784

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993N IN UGC 5695 J. Mueller reports her discovery of an apparent supernova located 6′′.7

west and 9′′.4 north of the nucleus of UGC 5695 (α = 10^{h}27^{m}.1, δ = +13^{o}16′, equinox 1950.0). The new object was found at mag ∼ 17.5 on a blue plate taken Apr. 15 by C. Brewer and Mueller with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey.

E. Cappellaro and M. Della Valle, European Southern Observatory, communicate: “CCD frames obtained on May 6.1 UT at the 1.5-m Danish telescope at La Silla confirm the presence of the new object 6′′.8 west and 9′′.8 north of the nucleus of UGC 5695. Preliminary photometry gives: V = 18.6, B–V = +0.5, V–R = +0.7.”

SUPERNOVA 1993J, 3C 273, AND 3C 279 E. I. Robson, Joint Astronomy Centre, Hilo, reports: “The James

Clerk Maxwell Telescope has been observing SN 1993J since outburst at a wavelength of 1.1 mm using the position given on IAUC 5752. No detections have been obtained; the lowest level reached was < 13.6 mJy (3σ) on Apr. 18 UT. This was in general agreement with the latest OVRO data on IAUC 5775, showing the decline in flux to higher radio frequencies, and with the IRAM data of IAUC 5768. All observing conditions have been excellent, and calibration was done using Mars. In terms of flux calibration, we note that the blazars 3C 273 and 3C 279 are both currently flaring; 3C 273 has increased by a factor of 2 at 1.1 mm since the beginning of 1993, 3C 279 by 1.5 over the same period.”

NOVA OPHIUCHI 1993 Further photometry: Apr. 24.63 UT, V = 9.82, B–V = +1.49 (W.

Allen, Adams Lane Observatory; cf. IAUC 5774); May 1.666, V = 11.85 ± 0.02, U–B = –0.19 ± 0.02, B–V = +0.84 ± 0.00, V–R = +1.89 ± 0.03, V–I = +1.83 ± 0.06 (A. C. Gilmore, Mt. John University Observatory).

BL LACERTAE The following visual magnitude estimates have been reported (cf. IAUC

5765): 1992 Dec. 29.70 UT, 15.6 (L. Szentaskó, Budapest, Hungary); 1993 Jan. 17.76, 15.1 (M. Verdenet, Bourbon Lancy, France); Apr. 20.12, 13.5 (Verdenet); 24.11, 13.8 (Verdenet).

1993 May 6 Daniel W. E. Green